In the aggregate expenditure model for a closed economy, assuming investment, government spending and taxes are exogenous, if the marginal propensity to consume is 0.8, a simultaneous 50 unit increase in government spending and a 20 unit decrease in investment will change equilibrium income by:


350 units.


87.5 units.


150 units.


500 units.


1
Expert's answer
2021-09-14T07:30:29-0400

Solution:

The correct answer is 150 units.

Equilibrium income will change by 150 units.

 

Given MPC of 0.8, derive the spending multiplier:

Multiplier = "\frac{1}{1 - MPC} = \frac{1}{1 - 0.8} = \frac{1}{0.2} = 5"


Multiplier = 5


"\triangle Y = (\triangle I\times M)+ (\triangle G \times M)"


"\triangle Y = (-20\times 5)+ (50 \times M)"


ΔY = (-20 x 5) + (50 x 5) = -100 + 250 = 150 units